DragonSlaYeR:

Evaluate the expression: 5y - 3 when y = 4 Evaluate the expression: -3r - 4 when r = 7 Evaluate the expression: (4w−3)2(4w−3)2 when w = 2

DragonSlaYeR:

@dontsaymyname

DragonSlaYeR:

theres three

DragonSlaYeR:

*there's

dontsaymyname:

Alrighty sooo all we're doing is plugging in the numbers just like last time :) 1. 5(4) - 3 2. -3(7) - 4 3. (4(2)−3)2(4(2)−3)2 Let me know what answers you get :)

DragonSlaYeR:

1. 17?

DragonSlaYeR:

2. -25

DragonSlaYeR:

dk 3.

dontsaymyname:

1 and 2 are correct

DragonSlaYeR:

dk 3?

dontsaymyname:

I'll step you three 1 sec

dontsaymyname:

((4)(2)−3)(2)((4)(2)−3)(2) =(8−3)(2)((4)(2)−3)(2) =(5)(2)((4)(2)−3)(2) =10((4)(2)−3)(2) =10(8−3)(2) =(10)(5)(2) Can you solve from here?

DragonSlaYeR:

100?

dontsaymyname:

Correct! Great job :)

DragonSlaYeR:

it's not

DragonSlaYeR:

it's 25

dontsaymyname:

Yeah you accidently put the equation twice